# GRADEO: Towards Human-Like Evaluation for Text-to-Video Generation via Multi-Step Reasoning

### v1

@@ Abstract
Recent great advances in video generation models have demonstrated their potential to produce high-quality videos, bringing challenges to effective evaluation. Unlike human evaluation, existing automated evaluation metrics lack high-level semantic understanding and reasoning capabilities for video, thus making them infeasible and unexplainable. To fill this gap, we curate GRADEO-Instruct, a multi-dimensional T2V evaluation instruction tuning dataset, including 3.3k videos from over 10 existing video generation models and multi-step reasoning assessments converted by 16k human annotations. We then introduce GRADEO, one of the first specifically designed video evaluation models, which grades AI-generated videos for explainable scores and assessments through multi-step reasoning. Experiments show that our method aligns better with human evaluations than existing methods. Furthermore, our benchmarking reveals that current video generation models struggle to produce content that aligns with human reasoning and complex real-world scenarios. The models, datasets, and codes will be released soon.
@@ 1 Introduction
Recent breakthroughs in Text-to-Video (T2V) generation models (Zhang et al., 2024a; Guo et al., 2024; Wang et al., 2024b; Zheng et al., 2024; OpenAI, 2024) have led to remarkable advancements in the field, demonstrating the potential to generate long-duration and realistic videos. However, existing models still generate hallucinated content, videos that do not align with real-world scenarios, or completely wrong videos. (Huang et al., 2024; Liu et al., 2024b, c; Kou et al., 2024; He et al., 2024; Bansal et al., 2024) Thus, there is a significant demand for automated frameworks that can objectively, accurately, and efficiently evaluate T2V models across diverse test sets, facilitating continuous performance and quality improvements.
@@
Early T2V generation evaluation method (Chivileva et al., 2023) introduces the T2V-CL framework, which focuses on image naturalness and text similarity but oversimplifies the evaluation. It also highlights the limitations of commonly used video quality metrics, including FVD (Unterthiner et al., 2018), CLIPSim (Wu et al., 2021), and IS (Salimans et al., 2016), and emphasizes that these metrics fail to eliminate the need for human assessment. While human preference remains the gold standard, it is not scalable to meet the growing demand for AI-generated video content. To automatically and comprehensively evaluate AI-generated videos, EvalCrafter (Liu et al., 2024b) and Vbench (Huang et al., 2024) propose detailed evaluation frameworks with hierarchical dimensions and automatic metrics, validated by human evaluations. On the other hand, some studies (Wu et al., 2024; Zhang et al., 2024d; Qu et al., 2024) attempt to model human evaluation using large-scale datasets that mimic human scores, aiming for a unified and simplified evaluation method.
Nevertheless, critical challenges persist in the evaluation of text-to-video generation: First, metrics in multi-dimensional evaluation frameworks often rely on pre-trained models (Caron et al., 2021; Radford et al., 2021; Wu et al., 2023a) trained on large-scale real-world image and video datasets. While these models capture general visual features and reflect human preferences for real-world content, they struggle to represent the distributions of generated videos, limiting their ability to assess generated content accurately. Second, traditional non-LLMs evaluation methods, constrained by limited datasets and model parameters, lack the capability to accurately comprehend video content. As a result, they primarily focus on low-level perceptual aspects like visual quality and fail to evaluate high-level semantic dimensions, such as toxicity, alignment with real-world scenarios, and creativity. Several research (Bansal et al., 2024; Meng et al., 2024; Miao et al., 2024) focus on evaluating the physical commonsense or safety, but their methods are not scalable and lack comprehensive evaluation. Third, existing metrics, including SOTA method VideoScore (Kou et al., 2024; Lin et al., 2024b; He et al., 2024), provide only scores (cf. Figure 1) without offering explanations for the rationale behind the scoring, resulting in poor interpretability. Fourth, Multimodal Large Language Models (MLLMs) (OpenAI, 2023; Liu et al., 2023; Team et al., 2023), despite their potential to comprehend video content and evaluate videos based on given instructions, are not specifically trained for video evaluation and lack alignment with human preferences. As a result, they struggle to provide accurate and comprehensive assessments of generated videos.
@@ 2 Related Work
In recent years, diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2021; Ho et al., 2020) have emerged as the primary methods for text-to-video generation. The early video diffusion model VDM (Ho et al., 2022b) extends the architecture to a 3D U-Net, enabling joint training on both image and video data. Imagen Video (Ho et al., 2022a), Video LDM (Blattmann et al., 2023), LaVie (Wang et al., 2024b), and Show-1 (Zhang et al., 2024a) leverage multi-stage models for high-resolution video generation. Some training-free methods (Khachatryan et al., 2023; Zhang et al., 2024c) modify only the inference process to generate videos but face issues with frame inconsistency. AnimateDiff (Guo et al., 2024) trains a plug-and-play motion module that can be integrated into any text-to-image model, enabling broad community applications. After the emergence of Sora (OpenAI, 2024), many models pre-trained on large-scale video datasets have been made open-source, including Open-sora (Zheng et al., 2024), Open-sora-plan (Lin et al., 2024a), and Mochi-1 (Team, 2024). Despite their impressive generation performance, closed-source commercial models (OpenAI, 2024; Runway, 2024; Lab, 2024; Kling, 2024; Zhipu, 2024) exhibit significantly stronger performance. However, these models fail to align with human evaluations in areas such as simulating real-world scenarios and ensuring safety. To address these limitations, we incorporate multimodal reasoning large models to assess these dimensions more effectively.
@@
Built upon Large Language Models (LLMs) (OpenAI, 2022; Achiam et al., 2023; Bai et al., 2023; Touvron et al., 2023a, b) and vision encoders (Radford et al., 2021; He et al., 2022), Multimodal Large Language Models (MLLMs) (OpenAI, 2023; Liu et al., 2023; Team et al., 2023) exhibit artificial general intelligence, excelling in open-world vision tasks such as understanding, instruction following, and conversational abilities. In-Context Learning (ICL) (Brown et al., 2020) enable LLMs excellent reasoning capability by providing samples and context. Additionally, Chain-of-Thought (CoT) (Wei et al., 2022) further strengthens multi-step reasoning by generating intermediate reasoning steps to solve complex problems. Extensions of CoT, such as Multimodal-CoT (Zhang et al., 2023), VCoT (Rose et al., 2023), and Visual CoT (Shao et al., 2024), explore visual reasoning in Multimodal Large Language Models (MLLMs). Recent works (Hao et al., 2024; Wang et al., 2024c) apply CoT to the video understanding domain. LLava-CoT and Marco-o1 (Xu et al., 2024; Zhao et al., 2024) integrate CoT into the model training process, utilizing organized tags to enable a systematic and structured reasoning process. Inspired by these work, we aim to explore the potential of reasoning MLLMs in text-to-video evaluation.
@@
With text-to-video models show exceptional performance on high resolution video generation, there is increasing interest in evaluating text-to-video models. Fréchet Video Distance (FVD) (Unterthiner et al., 2018), Inception Score (IS) (Salimans et al., 2016) and CLIP-Similarity (Radford et al., 2021) are most commonly automatic video quality assessment metrics. However, there are concerns about correlation between them and the evaluation of humans. VBench (Huang et al., 2024) introduces multi-dimensional and multi-categorical benchmark suite for evaluating video generation model performance. EvalCrafter (Liu et al., 2024b) construct their metrics by human alignment. Some recent studies(Bansal et al., 2024; Sun et al., 2024; Meng et al., 2024) prompt MLLMs to assess video quality or other performance dimension of generative models. VideoScore (He et al., 2024) shows models that were not trained on video evaluation task did not correlate well with human assessments. We propose a comprehensive benchmark suite including seven dimensions from video low-level perception to high-level video semantic understanding and common sense reasoning. Also, we introduce an automatic model for fine-grained evaluating text-to-video models.
@@ 3.1 GRADEO Evaluation Dimensions
Existing methods (Liu et al., 2024b; Huang et al., 2024; Kou et al., 2024) typically focus on assessing low-level dimensions, primarily video quality, with relatively less attention given to higher-level semantic evaluations. In addition to retaining the primary dimensions of assessment from previous methods, we include evaluations on whether the video complies with human common sense, whether it is safe, and whether it possesses creativity. As a result, we develop a bottom-up benchmark that spans from low-level visual perception to high-level visual understanding and reasoning. As shown in Table 1, we also define key aspects for each dimension with fine granularity. Further details are provided in Appendix A.1.
@@ 3.2 GRADEO-Instruct Dataset Construction
To construct a comprehensive prompt dataset for text-to-video evaluation, we ensured broad coverage across the dimensions of Quality, Aesthetics, Consistency, and Alignment. This was achieved by collecting and sampling from large-scale text-to-video pretraining dataset WebVid (Bain et al., 2021), real user-generated prompt dataset VidProM (Wang & Yang, 2024), and existing open-source text-to-video benchmark datasets (Liu et al., 2024b; Huang et al., 2024; Sun et al., 2024; Feng et al., 2024; Liu et al., 2024c; Kou et al., 2024). Additionally, the Safety dimension prompts are curated from specialized safety-focused dataset SafeSora (Dai et al., 2024). Inspired by recent studies (Bansal et al., 2024; Fu et al., 2024; Meng et al., 2024) on evaluating text-to-visual generation based on physical laws, we define thematic categories (e.g., Physical Laws, Emotional Responses, Weather Conditions, Chemical Laws, see more themes and prompt examples in Appendix A.2) that encompass human commonsense knowledge and manually craft multiple examples for each category in Rationality dimension. These examples guide the GPT-assisted generation of additional prompts. For Creativity dimension, we collect prompts from the video generation community to ensure diverse and imaginative inputs, while also generating additional creative prompts based on community examples. At the same time, we include simpler, less creative prompts (e.g., a bird flying in the sky, a child playing with a ball) to maintain a balanced dataset.
@@
For the dimensions of Quality, Aesthetics, Consistency, Alignment, and Rationality, we leverage a range of open-source models, including Open-Sora (Zheng et al., 2024), VideoCrafter-1 (Chen et al., 2023), VideoCrafter-2 (Chen et al., 2024a), Latte (Ma et al., 2024), LaVie (Wang et al., 2024b), and ZeroScope-576w (Cerspense, 2024), to generate videos. Moreover, we collect pre-generated videos from the VidProM dataset, ensuring a diverse set of examples. For the Safety dimension, we sample malicious videos from the SafeSora dataset and integrate benign videos from VidProM. For the Creativity dimension, we use the Kling 1.0 and Kling 1.5 models (Kling, 2024) to generate videos with varying creativity levels, enhancing the diversity of the dataset. Finally, We filter out videos with unsuitable dynamics to ensure they meet the required motion quality, and crop those with watermarks to eliminate model-specific characteristics. Details are shown in Appendix A.2.
@@
To create an instruction tuning dataset that enables multimodal large models to learn human evaluation of AI-generated videos, we define an assessing Chain-of-Thought process with four steps: <Overview>, <Description>, <Analysis>, and <Assessment>. The overview step involves planning the overall reasoning process for the subsequent stages. The description step provides detailed descriptions of the video content. The analysis step emulates the slow thinking and reasoning process in human evaluation, systematically elaborating on the rationale behind the scores. Finally, the assessment step generates the evaluation score based on the reasoning and analysis. We designed a GPT-assisted CoT response construction pipeline, where we feed three keyframes from the video, evaluation criteria, and human annotators’ assessment rationales into GPT-4o (Hurst et al., 2024) to generate CoT data consisting of four steps. If the generated score in the evaluation step does not align with the human-assigned score, we discard the CoT data and regenerate it. For chains with correct scores, we manually verified that the reasoning steps were consistent with human logic.
@@ 4 Experiments
$\rho_{srocc}$($\uparrow$)/$\rho_{plcc}$($\uparrow$)/$\bar{\Delta}$($\downarrow$) $\rho_{srocc}$($\uparrow$)/$\rho_{plcc}$($\uparrow$)/$\bar{\Delta}$($\downarrow$) $\rho_{srocc}$($\uparrow$)/$\rho_{plcc}$($\uparrow$)/$\bar{\Delta}$($\downarrow$) $\rho_{srocc}$($\uparrow$)/$\rho_{plcc}$($\uparrow$)/$\bar{\Delta}$($\downarrow$) $\rho_{srocc}$($\uparrow$)/$\rho_{plcc}$($\uparrow$)/$\bar{\Delta}$($\downarrow$) $\rho_{srocc}$($\uparrow$)/$\rho_{plcc}$($\uparrow$)/$\bar{\Delta}$($\downarrow$) $\rho_{srocc}$($\uparrow$)/$\rho_{plcc}$($\uparrow$)/$\bar{\Delta}$($\downarrow$)
@@
PIQE (N et al., 2015) 0.053/-0.040/1.123 -0.071/-0.259/1.123 0.045/-0.142/ 1.545 0.145/0.004/1.544 -0.124/-0.254/1.460 0.078/-0.154/1.440 0.361/0.305/1.208
BRISQUE (Mittal et al., 2012) 0.079/-0.049/1.193 0.021/-0.327/1.509 0.019/-0.338/1.750 0.085/-0.085/1.772 -0.063/-0.188/1.240 0.234/0.009/0.720 0.335/0.029/1.250
CLIP-Score (Hessel et al., 2021) 0.180/0.088/1.070 0.464/0.377/1.211 0.276/0.177/1.432 0.201/0.051/1.404 0.181/0.063/1.160 0.331/0.147/1.080 0.396/0.342/1.000
@@
LLaVA-1.5-7B (Liu et al., 2023) 0.563/0.176/0.684 0.544/0.199/0.518 0.273/0.194/0.705 0.405/-0.256/0.589 0.273/0.137/1.041 0.493/0.280/1.400 0.607/0.566/0.958
LLaVA-1.6-7B (Liu et al., 2024a) 0.514/0.132/0.667 0.567/0.434/0.500 0.334/0.103/0.591 0.421/0.182/2.036 0.137/-0.096/3.612 0.663/0.568/1.360 0.700/0.581/0.875
GLM-4v (GLM et al., 2024) 0.543/0.429/0.895 0.412/0.305/0.964 0.381/0.237/1.068 0.333/0.070/0.893 0.580/0.480/1.080 0.584/0.496/1.280 0.445/0.380/1.292
@@
InternVL2.5-8B (Chen et al., 2024b) 0.242/0.145/1.105 0.393/0.302/0.911 0.192/-0.061/1.091 0.194/0.030/2.054 0.322/0.306/1.255 0.563/0.547/1.080 0.528/0.381/1.208
LLaVA-NeXT-Video-7B (Zhang et al., 2024b) 0.585/0.441/0.825 0.260/0.036/0.930 0.290/-0.012/0.795 0.070/-0.127/0.895 0.350/0.137/1.204 0.602/0.038/1.900 0.523/0.291/1.167
Gemini-1.5-Flash (Team et al., 2023) 0.617/0.510/0.825 0.663/0.604/0.518 0.322/0.000/0.791 0.163/0.048/1.446 0.071/-0.033/1.449 0.348/0.255/1.380 0.026/-0.035/1.375
Gemini-1.5-Pro (Team et al., 2023) 0.538/0.454/0.614 0.589/0.556/0.526 0.401/0.238/0.886 0.488/0.283/1.161 0.288/0.227/1.469 0.619/0.44/1.700 0.792/0.721/0.792
VideoScore-v1.1 (He et al., 2024) 0.365/-0.108/0.789 - 0.575/0.386/0.545 0.484/0.287/0.649 0.297/0.070/0.780 - -

### v2

@@ Abstract
Recent great advances in video generation models have demonstrated their potential to produce high-quality videos, bringing challenges to effective evaluation. Unlike human evaluation, existing automated evaluation metrics lack high-level semantic understanding and reasoning capabilities for video, thus making them infeasible and unexplainable. To fill this gap, we curate GRADEO-Instruct, a multi-dimensional T2V evaluation instruction tuning dataset, including 3.3k videos from over 10 existing video generation models and multi-step reasoning assessments converted by 16k human annotations. We then introduce GRADEO, one of the first specifically designed video evaluation models, which grades AI-generated videos for explainable scores and assessments through multi-step reasoning. Experiments show that our method aligns better with human evaluations than existing methods. Furthermore, our benchmarking reveals that current video generation models struggle to produce content that aligns with human reasoning and complex real-world scenarios.
@@ 1 Introduction
Recent breakthroughs in Text-to-Video (T2V) generation models (zhang2023show; guo2024animatediff; wang2023lavie; opensora; sora) have led to remarkable advancements in the field, demonstrating the potential to generate long-duration and realistic videos. However, existing models still generate hallucinated content, videos that do not align with real-world scenarios, or completely wrong videos. (huang2024vbench; liu2024evalcrafter; liu2024fetv; kou2024subjective; he2024videoscore; bansal2024videophy) Thus, there is a significant demand for automated frameworks that can objectively, accurately, and efficiently evaluate T2V models across diverse test sets, facilitating continuous performance and quality improvements.
@@
Early T2V generation evaluation method (chivileva2023measuring) introduces the T2V-CL framework, which focuses on image naturalness and text similarity but oversimplifies the evaluation. It also highlights the limitations of commonly used video quality metrics, including FVD (unterthiner2018towards), CLIPSim (wu2021godiva), and IS (salimans2016improved), and emphasizes that these metrics fail to eliminate the need for human assessment. While human preference remains the gold standard, it is not scalable to meet the growing demand for AI-generated video content. To automatically and comprehensively evaluate AI-generated videos, EvalCrafter (liu2024evalcrafter) and Vbench (huang2024vbench) propose detailed evaluation frameworks with hierarchical dimensions and automatic metrics, validated by human evaluations. On the other hand, some studies (wu2024towards; zhang2024benchmarkingmultidimensionalaigcvideo; qu2024exploring) attempt to model human evaluation using large-scale datasets that mimic human scores, aiming for a unified and simplified evaluation method.
Nevertheless, critical challenges persist in the evaluation of text-to-video generation: First, metrics in multi-dimensional evaluation frameworks often rely on pre-trained models (Caron_2021_ICCV; radford2021learning; wu2023dover) trained on large-scale real-world image and video datasets. While these models capture general visual features and reflect human preferences for real-world content, they struggle to represent the distributions of generated videos, limiting their ability to assess generated content accurately. Second, traditional non-LLMs evaluation methods, constrained by limited datasets and model parameters, lack the capability to accurately comprehend video content. As a result, they primarily focus on low-level perceptual aspects like visual quality and fail to evaluate high-level semantic dimensions, such as toxicity, alignment with real-world scenarios, and creativity. Several research (bansal2024videophy; meng2024towards; miao2024t2vsafetybench) focus on evaluating the physical commonsense or safety, but their methods are not scalable and lack comprehensive evaluation. Third, existing metrics, including SOTA method VideoScore (kou2024subjective; lin2024evaluating; he2024videoscore), provide only scores (cf. Figure 1) without offering explanations for the rationale behind the scoring, resulting in poor interpretability. Fourth, Multimodal Large Language Models (MLLMs) (gpt4v; llava; team2023gemini), despite their potential to comprehend video content and evaluate videos based on given instructions, are not specifically trained for video evaluation and lack alignment with human preferences. As a result, they struggle to provide accurate and comprehensive assessments of generated videos.
@@ 2 Related Work
In recent years, diffusion models (sohl2015deep; song2021scorebased; ho2020denoising) have emerged as the primary methods for text-to-video generation. The early video diffusion model VDM (ho2022video) extends the architecture to a 3D U-Net, enabling joint training on both image and video data. Imagen Video (ho2022imagen), Video LDM (blattmann2023align), LaVie (wang2023lavie), and Show-1 (zhang2023show) leverage multi-stage models for high-resolution video generation. Some training-free methods (khachatryan2023text2video; zhang2024controlvideo) modify only the inference process to generate videos but face issues with frame inconsistency. AnimateDiff (guo2024animatediff) trains a plug-and-play motion module that can be integrated into any text-to-image model, enabling broad community applications. After the emergence of Sora (sora), many models pre-trained on large-scale video datasets have been made open-source, including Open-sora (opensora), Open-sora-plan (lin2024open), and Mochi-1 (genmo2024mochi). Despite their impressive generation performance, closed-source commercial models (sora; runwaygen2; pika; kling; qingying) exhibit significantly stronger performance. However, these models fail to align with human evaluations in areas such as simulating real-world scenarios and ensuring safety. To address these limitations, we incorporate multimodal reasoning large models to assess these dimensions more effectively.
@@
Built upon Large Language Models (LLMs) (gpt3.5; gpt4; bai2023qwen; llama1; llama2) and vision encoders (radford2021learning; he2022masked), Multimodal Large Language Models (MLLMs) (gpt4v; llava; team2023gemini) exhibit artificial general intelligence, excelling in open-world vision tasks such as understanding, instruction following, and conversational abilities. In-Context Learning (ICL) (gpt3) enable LLMs excellent reasoning capability by providing samples and context. Additionally, Chain-of-Thought (CoT) (wei2022chain) further strengthens multi-step reasoning by generating intermediate reasoning steps to solve complex problems. Extensions of CoT, such as Multimodal-CoT (zhang2023multimodal), VCoT (rose2023visual), and Visual CoT (shao2024visual), explore visual reasoning in Multimodal Large Language Models (MLLMs). Recent works (vot; wang2024videocot) apply CoT to the video understanding domain. LLava-CoT and Marco-o1 (xu2024llava; zhao2024marco) integrate CoT into the model training process, utilizing organized tags to enable a systematic and structured reasoning process. Inspired by these work, we aim to explore the potential of reasoning MLLMs in text-to-video evaluation.
@@
With text-to-video models show exceptional performance on high resolution video generation, there is increasing interest in evaluating text-to-video models. Fréchet Video Distance (FVD) (unterthiner2018towards), Inception Score (IS) (salimans2016improved) and CLIP-Similarity (radford2021learning) are most commonly automatic video quality assessment metrics. However, there are concerns about correlation between them and the evaluation of humans. VBench (huang2024vbench) introduces multi-dimensional and multi-categorical benchmark suite for evaluating video generation model performance. EvalCrafter (liu2024evalcrafter) construct their metrics by human alignment. Some recent studies(bansal2024videophy; sun2024t2v; meng2024towards) prompt MLLMs to assess video quality or other performance dimension of generative models. VideoScore (he2024videoscore) shows models that were not trained on video evaluation task did not correlate well with human assessments. Evaluation Agent (zhang2024evaluationagent) introduces a dynamic, agent-based framework with hierarchical assessments but raises fairness concerns in evaluating diverse models while it lacks human annotations. We propose a comprehensive benchmark suite including seven dimensions from video low-level perception to high-level video semantic understanding and common sense reasoning. Also, we introduce an automatic model for fine-grained evaluating text-to-video models.
@@ 3.1 GRADEO Evaluation Dimensions
Existing methods (liu2024evalcrafter; huang2024vbench; kou2024subjective) typically focus on assessing low-level dimensions, primarily video quality, with relatively less attention given to higher-level semantic evaluations. In addition to retaining the primary dimensions of assessment from previous methods, we include evaluations on whether the video complies with human common sense, whether it is safe, and whether it possesses creativity. As a result, we develop a bottom-up benchmark that spans from low-level visual perception to high-level visual understanding and reasoning. As shown in Table 1, we also define key aspects for each dimension with fine granularity. Further details are provided in Appendix A.1.
@@ 3.2 GRADEO-Instruct Dataset Construction
To construct a comprehensive prompt dataset for text-to-video evaluation, we ensured broad coverage across the dimensions of Quality, Aesthetics, Consistency, and Alignment. This was achieved by collecting and sampling from large-scale text-to-video pretraining dataset WebVid (Bain21), real user-generated prompt dataset VidProM (wang2024vidprom), and existing open-source text-to-video benchmark datasets (liu2024evalcrafter; huang2024vbench; sun2024t2v; feng2024tc; liu2024fetv; kou2024subjective). Additionally, the Safety dimension prompts are curated from specialized safety-focused dataset SafeSora (dai2024safesora). Inspired by recent studies (bansal2024videophy; fu2024commonsense; meng2024towards) on evaluating text-to-visual generation based on physical laws, we define thematic categories (e.g., Physical Laws, Emotional Responses, Weather Conditions, Chemical Laws, see more themes and prompt examples in Appendix A.2) that encompass human commonsense knowledge and manually craft multiple examples for each category in Rationality dimension. These examples guide the GPT-assisted generation of additional prompts. For Creativity dimension, we collect prompts from the video generation community to ensure diverse and imaginative inputs, while also generating additional creative prompts based on community examples. At the same time, we include simpler, less creative prompts (e.g., a bird flying in the sky, a child playing with a ball) to maintain a balanced dataset.
@@
For the dimensions of Quality, Aesthetics, Consistency, Alignment, and Rationality, we leverage a range of open-source models, including Open-Sora (opensora), VideoCrafter-1 (chen2023videocrafter1), VideoCrafter-2 (chen2024videocrafter2), Latte (ma2024latte), LaVie (wang2023lavie), and ZeroScope-576w (zeroscope), to generate videos. Moreover, we collect pre-generated videos from the VidProM dataset, ensuring a diverse set of examples. For the Safety dimension, we sample malicious videos from the SafeSora dataset and integrate benign videos from VidProM. For the Creativity dimension, we use the Kling 1.0 and Kling 1.5 models (kling) to generate videos with varying creativity levels, enhancing the diversity of the dataset. Finally, We filter out videos with unsuitable dynamics to ensure they meet the required motion quality, and crop those with watermarks to eliminate model-specific characteristics. Details are shown in Appendix A.2.
@@
To create an instruction tuning dataset that enables multimodal large models to learn human evaluation of AI-generated videos, we define an assessing Chain-of-Thought process with four steps: <Overview>, <Description>, <Analysis>, and <Assessment>. The overview step involves planning the overall reasoning process for the subsequent stages. The description step provides detailed descriptions of the video content. The analysis step emulates the slow thinking and reasoning process in human evaluation, systematically elaborating on the rationale behind the scores. Finally, the assessment step generates the evaluation score based on the reasoning and analysis. We designed a GPT-assisted CoT response construction pipeline, where we feed three keyframes from the video, evaluation criteria, and human annotators’ assessment rationales into GPT-4o (hurst2024gpt) to generate CoT data consisting of four steps. If the generated score in the evaluation step does not align with the human-assigned score, we discard the CoT data and regenerate it. For chains with correct scores, we manually verified that the reasoning steps were consistent with human logic.
@@ 4 Experiments
[…]
@@
PIQE (piqe) 0.053/-0.040/1.123 -0.071/-0.259/1.123 0.045/-0.142/ 1.545 0.145/0.004/1.544 -0.124/-0.254/1.460 0.078/-0.154/1.440 0.361/0.305/1.208
BRISQUE (BRISQUE) 0.079/-0.049/1.193 0.021/-0.327/1.509 0.019/-0.338/1.750 0.085/-0.085/1.772 -0.063/-0.188/1.240 0.234/0.009/0.720 0.335/0.029/1.250
CLIP-Score (hessel2021clipscore) 0.180/0.088/1.070 0.464/0.377/1.211 0.276/0.177/1.432 0.201/0.051/1.404 0.181/0.063/1.160 0.331/0.147/1.080 0.396/0.342/1.000
@@
LLaVA-1.5-7B (llava) 0.563/0.176/0.684 0.544/0.199/0.518 0.273/0.194/0.705 0.405/-0.256/0.589 0.273/0.137/1.041 0.493/0.280/1.400 0.607/0.566/0.958
LLaVA-1.6-7B (liu2024llavanext) 0.514/0.132/0.667 0.567/0.434/0.500 0.334/0.103/0.591 0.421/0.182/2.036 0.137/-0.096/3.612 0.663/0.568/1.360 0.700/0.581/0.875
GLM-4v (glm2024chatglm) 0.543/0.429/0.895 0.412/0.305/0.964 0.381/0.237/1.068 0.333/0.070/0.893 0.580/0.480/1.080 0.584/0.496/1.280 0.445/0.380/1.292
@@
InternVL2.5-8B (chen2024internvl) 0.242/0.145/1.105 0.393/0.302/0.911 0.192/-0.061/1.091 0.194/0.030/2.054 0.322/0.306/1.255 0.563/0.547/1.080 0.528/0.381/1.208
LLaVA-NeXT-Video-7B (zhang2024llavanext-video) 0.585/0.441/0.825 0.260/0.036/0.930 0.290/-0.012/0.795 0.070/-0.127/0.895 0.350/0.137/1.204 0.602/0.038/1.900 0.523/0.291/1.167
Qwen2-VL-7B (Qwen2VL) 0.590/0.424/0.649 0.648/0.599/0.667 0.491/0.310/0.682 0.233/-0.022/1.000 0.461/0.225/1.060 0.559/0.244/2.320 0.606/0.593/1.083
Gemini-1.5-Flash (team2023gemini) 0.617/0.510/0.825 0.663/0.604/0.518 0.322/0.000/0.791 0.163/0.048/1.446 0.071/-0.033/1.449 0.348/0.255/1.380 0.026/-0.035/1.375
Gemini-1.5-Pro (team2023gemini) 0.538/0.454/0.614 0.589/0.556/0.526 0.401/0.238/0.886 0.488/0.283/1.161 0.288/0.227/1.469 0.619/0.44/1.700 0.792/0.721/0.792
VideoScore-v1.1 (he2024videoscore) 0.365/-0.108/0.789 - 0.575/0.386/0.545 0.484/0.287/0.649 0.297/0.070/0.780 - -